OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find g(-3).

OpenStudy (loser66):

\[\huge f(x) =2x-4\\\huge g(x)= x^2 -4x\\find~g(-3)\] My question is: why do they give you f(x)?

OpenStudy (anonymous):

idk

OpenStudy (loser66):

is there any question else?

OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find f(0).

OpenStudy (anonymous):

i think when it says f(0), the 0 goes for x

OpenStudy (e.mccormick):

These work just like the other one Hero just showed you. Put what is inside the ( ) in place of x. Choose which function you put it on based on the letter next to ( ), so g(-3) would mean use function g and relace x with -3 and f(0) would mean sue function f and replace x with 0.

OpenStudy (anonymous):

ok